A cube has edges that are 6 centimeters long. How many times greater is the volume of a cube with 9 centimeter sides?
Answer by mananth(16946)

cube volume = 6*6*6=216 cu. cm
..
second cube volume = 9*9*9=729
..
729 /216 = 3.37 times bigger